OpenStudy (anonymous):

For which values of K satisfy the convergence of the method of Gauss-Jacobi in the following system

OpenStudy (anonymous):

the system................?

OpenStudy (anonymous):

\[\left\{\\ Kx + y + 2z = 2 \\ Ky - (K/5)z = 3 \\ Kx + y = 2 \\ \right\}\]

OpenStudy (anonymous):

i'm not sure but its value of K=1

OpenStudy (anonymous):

Please i need the steps of solving it...